Time value Personal Finance Problem You have S 1,500 to invest today at 7% interest compounded annually. a. Find how much you will have accumulated in the account at the end of (1) 3 years, (2) 6 years, and (3) 9 years. b. Use your findings in part a to calculate the amount of interest eamed in (1) the first 3 years (years 1 to 3), (2) the second 3 years (years 4 to 6), and (3) the third 3 years (years 7 to 9). c. Compare and contrast your findings in part b. Explain how the amount of interest earned changes in each succeeding 3-year peiod a. (1) Al the end of 3 years, the amount you will have accumulaled is (Round to the nearest cent) (2) At the end of 6 years, the amount you will have accumulated is SRound to the nearest cent.) (3) At the end of 9 years, the amount you will have accumulated is Round to the nearest cent.) b. (1) From your findings in part , the 8mount of interest earned in the first 3 years (years 1 to 3) is $ (Round to the nearest cent ) (2) From your findings in part a, the amount of interest earned in the second 3 years (years 4 to 6) is (Round to the nearest cent.) (3) From your findings in part a, the amount of interest earned in the third 3 years (years 7 to 9) is $ . (Round to the nearest cent) c. According to the findings in part b, the amount of inlerest earned in each succeeding 3-year period due to compournding Select from the drop-down menu)
Answer #1

a.

(1)

accumulated value=1500*(1+7%)^3=1837.56

(2)

accumulated value=1500*(1+7%)^6=2251.10

(3)

accumulated value=1500*(1+7%)^9=2757.69

b.

(1)

interest earned=1837.56-1500=337.56

(2)

interest earned=2251.10-1837.56=413.54

(3) increases

the above is answer..

interest earned=2757.69-2251.10=506.59

c.